# Efficient construction of the Feynman-Vernon influence functional as matrix product states

### v1

@@ I Introduction
For small impurities, such as a single spin or a single electron, the construction of the MPS representation of the Feynman-Vernon IF (referred to as the MPS-IF afterwards) is the dominant calculation in TEMPO. The approach originally taken in TEMPO [33], is to build up the MPS-IF by decomposing the IF into the product of many partial IFs, each written as an MPS of a small bond dimension and the total number of partial IFs scales linearly with the total evolution time. Another approach to build the MPS-IF, which is used in the tensor network IF methods [47, 48], is to map the IF into a Gaussian state and then construct it by applying an equivalent quantum circuit of local gate operations onto the vacuum state using the Fisherman-White (FW) algorithm [58]. The depth of the quantum circuit has been shown to scale only logarithmically with the evolution time in certain cases. The partial IF approach has also been used in combination with the tensor network IF method, and is reported to have a similar accuracy with the FW algorithm [50]. In both approaches, the time-translationally invariant (TTI) property of the IF is explicitly destroyed.
@@ II.3 The time-translationally invariant IF method
where $\alpha_{l}$ and $\lambda_{l}$ are parameters to be determined, and $x$ can be both positive and negative. Finding the optimal $\alpha_{l}$ and $\lambda_{l}$ in Eq.(19) is an important and well-studied task in signal processing, which can be solved by the Prony algorithm [64]. We denote the error occurred in this decomposition as
@@ II.4 TTI approach to construct the MPS-IF for bosonic QIMs
where $\hat{H_{\rm S}}$ is the impurity spin Hamiltonian, $\hat{b}^{\dagger}_{k}$ and $\hat{b}_{k}$ are bosonic creation and annihilation operators.
@@ III.2 The single impurity Anderson model
As an application for a harder instance, we apply our method to study the steady state current of the single impurity Anderson model coupled to two baths. We use the same settings as used in Refs. [11, 48, 41], namely the two baths are both at zero temperature, with chemical potentials $\mu_{1}=-\mu_{2}=V/2$ and semi-circular bath spectrum density in Eq.(50) with $\Gamma=0.1$ (we use $\Gamma$ as the unit in this case). In Refs. [48, 41], the steady state particle current is computed by performing real-time evolution till $\Gamma t=4.2$. In particular, the results in Ref. [41] calculated by the partial IF method have well converged against different $\delta t$ and MPS bond truncation tolerance, therefore the major remaining factor that may affect the quality of the obtained steady state current is the total evolution time (the whole system may have not reached its non-equilibrium steady state yet with $\Gamma t=4.2$). With the more efficient TTI IF method to construct the MPS-IF, we can easily reach longer evolution time. In this work, we thus evolve the system till $\Gamma t=6.3$ and check if the previous results have well converged to their steady state values. We denote the particle current from the $\nu$th bath with spin $\sigma$ into the impurity as $\mathcal{J}^{\nu}_{\sigma}$ (see Ref. [41] for the definition of particle current and the way to calculate it based on the obtained MPS-IFs). As in Refs. [48, 41], we calculate the symmetric particle current $\mathcal{J}=(\mathcal{J}^{1}_{\uparrow}-\mathcal{J}^{2}_{\uparrow})/2=(\mathcal{J}^{1}_{\downarrow}-\mathcal{J}^{2}_{\downarrow})/2$.
@@
In Fig. 4, we plot the symmetric particle current versus time $t$, with the starting point $\Gamma t=4.2$, for $V$ from small to large. We have used $\Gamma\delta t=0.014$ and $\chi=160$ for all these simulations. As comparison, we have shown the partial IF results taken from Ref. [41] (under the same $\delta t$) with the same markers but in cyan (only for the starting time). The results from exact diagonalization (ED) is also shown in the same markers but in red (for the starting and final times). For ED results we have discretized each bath into $8000$ equal-distant frequencies and verified their convergence against bath discretization. We can see that the results calculated with our TTI IF method well matches with the previous partial IF results (the errors between them are within the first-order time discretization error). From Fig. 4(a, b, c) where $V/\Gamma<0.54$, we can see that the particle currents has converged fairly well for $U=0$, however, for $U>0$ the particle current increases significantly from $\Gamma t=4.2$ to $\Gamma t=6.3$, indicating that the whole system has not reached its steady state yet, and that the derivation from the steady state seems to be larger when $U/\Gamma$ increases from 00 to $4$. In comparison, from Fig. 4(d,e,f) where $V/\Gamma>0.7$, we can see that the particle currents have well converged for all values of $V$s and $U$s we have considered. Overall, the above results indicate that the system could reach its steady state more quickly for larger $V$ and smaller $U$.

### v2

@@ I Introduction
For small impurities, such as a single spin or a single electron, the construction of the MPS representation of the Feynman-Vernon IF (referred to as the MPS-IF afterwards) is the dominant calculation in TEMPO. The approach originally taken in TEMPO [33], is to build up the MPS-IF by decomposing the IF into the product of many partial IFs, each written as an MPS of a small bond dimension and the total number of partial IFs scales linearly with the total evolution time. Another approach to build the MPS-IF, which is used in the tensor network IF methods [47, 48], is to map the IF into a Gaussian state and then construct it by applying an equivalent quantum circuit of local gate operations onto the vacuum state using the Fishman-White (FW) algorithm [58]. The depth of the quantum circuit has been shown to scale only logarithmically with the evolution time in certain cases. The partial IF approach has also been used in combination with the tensor network IF method, and is reported to have a similar accuracy with the FW algorithm [50]. In both approaches, the time-translationally invariant (TTI) property of the IF is explicitly destroyed.
@@ II.3 The time-translationally invariant IF method
where $\alpha_{l}$ and $\lambda_{l}$ are parameters to be determined, and $x$ can be both positive and negative. Finding the optimal $\alpha_{l}$ and $\lambda_{l}$ in Eq.(19) is an important and well-studied task in signal processing, which can be solved by the Prony algorithm [64] (also see Appendix. A). We denote the error occurred in this decomposition as
@@ II.4 TTI approach to construct the MPS-IF for bosonic QIMs
where $\hat{H_{\rm S}}$ is the impurity spin Hamiltonian, $\hat{b}^{\dagger}_{k}$ and $\hat{b}_{k}$ are bosonic creation and annihilation operators. Here we also note that for bosonic impurity problems the TTI property of the IF has already been explored to speedup the construction of the MPS-IF [71, 72], however the strategies introduced in these works are very different from our approach.
@@ III.2 The single impurity Anderson model
As an application for a harder instance, we apply our method to study the steady state current of the single impurity Anderson model coupled to two baths. We use the same settings as used in Refs. [11, 48, 41], namely the two baths are both at zero temperature, with chemical potentials $\mu_{1}=-\mu_{2}=V/2$ and semi-circular bath spectrum density in Eq.(50) with $\Gamma=0.1$ (we use $\Gamma$ as the unit in this case). In Refs. [48, 41], the steady state particle current is computed by performing real-time evolution till $\Gamma t=4.2$. In particular, the results in Ref. [41] calculated by the partial IF method have well converged against different $\delta t$ and MPS bond truncation tolerance, therefore the major remaining factor that may affect the quality of the obtained steady state current is the total evolution time (the whole system may have not reached its non-equilibrium steady state yet with $\Gamma t=4.2$). With the more efficient TTI IF method to construct the MPS-IF, we can easily reach longer evolution time. In this work, we thus evolve the system till $\Gamma t=8.4$ and check if the previous results have well converged to their steady state values. We denote the particle current from the $\nu$th bath with spin $\sigma$ into the impurity as $\mathcal{J}^{\nu}_{\sigma}$ (see Ref. [41] for the definition of particle current and the way to calculate it based on the obtained MPS-IFs). As in Refs. [48, 41], we calculate the symmetric particle current $\mathcal{J}=(\mathcal{J}^{1}_{\uparrow}-\mathcal{J}^{2}_{\uparrow})/2=(\mathcal{J}^{1}_{\downarrow}-\mathcal{J}^{2}_{\downarrow})/2$.
@@
In Fig. 4, we plot the symmetric particle current versus time $t$, with the starting point $\Gamma t=4.2$, for $V$ from small to large. We have used $\Gamma\delta t=0.014$ and $\chi=160$ for all these simulations. As comparison, we have shown the partial IF results taken from Ref. [41] (under the same $\delta t$) with the same markers but in cyan (only for the starting time). The results from exact diagonalization (ED) is also shown in the same markers but in red (for the starting and final times). For ED results we have discretized each bath into $8000$ equal-distant frequencies and verified their convergence against bath discretization. We can see that the results calculated with our TTI IF method well matches with the previous partial IF results (the errors between them are within the first-order time discretization error). From Fig. 4(a, b, c) where $V/\Gamma<0.54$, we can see that the particle currents has converged fairly well for $U=0$, however, for $U>0$ the particle current increases significantly (especially for $U/\Gamma>2$) from $\Gamma t=4.2$ to $\Gamma t=8.4$, indicating that the whole system has not reached its steady state yet, and that the derivation from the steady state seems to be larger when $U/\Gamma$ increases from 00 to $4$. In comparison, from Fig. 4(d,e,f) where $V/\Gamma>0.7$, we can see that the particle currents have well converged for all values of $V$s and $U$s we have considered. Overall, the above results indicate that the system could reach its steady state more quickly for larger $V$ and smaller $U$.
@@ Acknowledgements.
Appendix A The Prony algorithm
We want to approximate a discrete function $\eta_{x}$ in the form (here for simplicity we assume $x\geq 0$)
$\eta_{x}\approx\sum_{l=1}^{n}\alpha_{l}\lambda_{l}^{x}.$ (52)
This is a nonlinear problem even if the value of $n$ is known. The goal is to pick $2n$ samples $\eta_{1},\ldots,\eta_{2n}$ to fit the $2n$ parameters $\alpha_{1},\ldots,\alpha_{n},\lambda_{1},\ldots,\lambda_{n}$. The above expression gives $n$ equations which may be expressed in the matrix form as
$\matrixquantity(\lambda_{1}^{0}&\cdots&\lambda_{n}^{0}\\ \vdots&\ddots&\vdots\\ \lambda_{1}^{n-1}&\cdots&\lambda_{n}^{n-1})\matrixquantity(\alpha_{1}\\ \vdots\\ \alpha_{n})=\matrixquantity(\eta_{1}\\ \vdots\\ \eta_{n}).$ (53)
Therefore if the $\lambda_{l}$s are determined, then we have a set of linear equations which can be easily solved to obtain the $\alpha_{l}$s.
The key of the Prony method is to recognize that Eq. (52) is the solution to some homogeneous linear constant-coefficient difference equation. In order to find such a difference equation, we define the characteristic polynomial $\phi(\lambda)$ as
$\phi(\lambda)=\prod_{l=1}^{n}(\lambda-\lambda_{l}).$ (54)
This polynomial has $\lambda_{l}$ as its roots and can be expaned into a power series as
$\phi(\lambda)=\sum_{k=0}^{n}a_{k}\lambda^{n-k},$ (55)
where the coefficient $a_{0}=1$. Employing Eq. (52), we have
$\begin{split}\sum_{k=0}^{n}a_{k}\eta_{p-k}=&\sum_{k=0}^{n}a_{k}\sum_{l=1}^{n}\alpha_{l}\lambda_{l}^{p-k-1}\\ =&\sum_{l=1}^{n}\alpha_{l}\lambda_{l}^{p-n-1}\sum_{k=0}^{n}a_{k}\lambda_{l}^{n-k}\\ =&\sum_{l=1}^{n}\alpha_{l}\lambda_{l}^{p-n-1}\phi(\lambda_{l})=0,\end{split}$ (56)
which is valid for $n+1\leq p\leq 2n$. Therefore we have $a_{0}\eta_{p}+\sum_{k=1}^{n}a_{k}\eta_{p-k}=0$, which can be expressed as an $n\times n$ matrix equation:
$\matrixquantity(\eta_{n}&\eta_{n-1}&\cdots&\eta_{1}\\ \eta_{n+1}&\eta_{n}&\cdots&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ \eta_{2n-1}&\eta_{2n-2}&\cdots&\eta_{n})\matrixquantity(a_{1}\\ a_{2}\\ \vdots\\ a_{n})=-\matrixquantity(\eta_{n+1}\\ \eta_{n+2}\\ \vdots\\ \eta_{2n}).$ (57)
Solving this matrix equation gives the coefficients $a_{k}$ of the characteristic polynomial $\phi(\lambda)$, and then its roots $\lambda_{l}$ can be calculated. Substituting $\lambda_{l}$ back into Eq. (53) and solving it to get the coefficients $\alpha_{l}$, and the desired expression (52) is obtained.